OpenStudy (anonymous):

find the derivative of f(x)=cosπx^2

myininaya (myininaya):

\[f(x)=\cos(\pi)x^2 or f(x)=\cos(\pi x^2) or f(x)=(\cos(\pi x))^2\] which is it?

OpenStudy (anonymous):

the second one I think because the book doesn't have any parenthesis at all

myininaya (myininaya):

\[f(x)=\cos(\pi x^2) => f'(x)=(\pi x^2)'[(\cos')(\pi x^2)]=\pi (x^2)'[(\cos')(\pi x^2)]\]

myininaya (myininaya):

\[f'(x)=\pi (2x)(-\sin(\pi x^2))\]

myininaya (myininaya):

\[f'(x)=-2\pi x \sin(\pi x^2)\]

OpenStudy (anonymous):

what rule did you use?

myininaya (myininaya):

chain rule do you not see i took the derivative of the inside and multiplied by derivative of outside?

myininaya (myininaya):

\[f(x)=h(g(x))=\cos(\pi x^2) => h(x)= \cos(x) , g(x)=\pi x^2\] \[f'(x)=g'(x)h'(g(x))=2\pi x(-\sin(g(x)))=2\pi x(-\sin(\pi x^2))=-2 \pi x \sin(\pi x^2)\]